Consider five proteins with the properties shown in the following table. Answer thequestions below about these proteins and justify your answers

  1. Protein

    Subunit mass (Da)

    Native mass (Da)

    A

    10,000

    40,000

    B

    15,000

    120,000

    C

    20,000

    20,000

    D

    25,000

    75,000

    E

    30,000

    60,000

    1. Which protein would migrate the slowest in an SDS-PAGE?

    2. Which protein would migrate the fast in a Native-PAGE?

    3. What is likely the quaternary structure (subunit composition) of each protein based on the data in the

      table?
Answer #1

Which protein will migrate slowest in SDS PAGE?

In denaturing SDS PAGE the proteins will be in monomer form. So that the subunit with largest molecular mass will migrate slower compared to the smaller ones. That is protein E with the highest mass (30kDa) will move slower.

Which protein would migrate the fast in a Native-PAGE?

In NATIVE PAGE protein will be in its normal form with the subunits attached to each other. So proteins with lower molecular weight (including the subunits) will be moving faster. Here protein C with native mass of 20kDa will move faster in the native –PAGE

Subunit composition of the proteins

We know that native mass in the sum of all the subunit of the protein so we can get a rough idea about the protein composition from the subunit mass and native mass by calculating the fraction from it.

Protein A- native mass 40000 Da and sub unit mass 10000 Da so it may containing 4 subunits and is a tetramer protein.

Like wise

Protein B- 120kDa/15kDA= 8 subunits

Protein C- has only one subunit because native mass= subunit mass

Protein D- 75kDA/25kDA= 3 subunits it is a trimer

Protein E- 60kDa/30kDa= 2 subunits and is a dimer.
Answer #2

1. Which protein would migrate the slowest in an SDS-PAGE?

  • SDS-PAGE separates proteins based on their subunit mass by denaturing them and giving them a uniform negative charge.

  • The slowest migrating protein would be the one with the largest subunit mass, which is:

    • Protein E (30,000 Da).

2. Which protein would migrate the fastest in a Native-PAGE?

  • Native-PAGE retains the native, folded structure and separates proteins based on size, shape, and charge.

  • The protein with the smallest native mass would migrate the fastest:

    • Protein C (20,000 Da).

3. What is likely the quaternary structure (subunit composition) of each protein based on the data in the table?

To determine the quaternary structure, divide the native mass by the subunit mass:

  • A:

40,00010,000=4    Tetramer\frac{40,000}{10,000} = 4 \implies \text{Tetramer}

  • B:

120,00015,000=8    Octamer\frac{120,000}{15,000} = 8 \implies \text{Octamer}

  • C:

20,00020,000=1    Monomer\frac{20,000}{20,000} = 1 \implies \text{Monomer}

  • D:

75,00025,000=3    Trimer\frac{75,000}{25,000} = 3 \implies \text{Trimer}

  • E:

60,00030,000=2    Dimer


answered by: anonymous